OpenStudy (anonymous):

Math Help Please! :) Evaluate f(x)= 3/5 x for x = 2.

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

@Mimi_x3

OpenStudy (anonymous):

@dumbcow

Nnesha (nnesha):

replace x by 2 and then solve

OpenStudy (haichi):

1.2

Nnesha (nnesha):

\[\huge\ f(2) =\frac{ 3 }{ 5 }x\]

Nnesha (nnesha):

replace x by 2

OpenStudy (anonymous):

F(2) = 3/5 (2)

Nnesha (nnesha):

yes right now just solve

OpenStudy (anonymous):

F (2) = 6/5 ?

OpenStudy (anonymous):

F (2) = 6/10

Nnesha (nnesha):

yes that's right :)

Nnesha (nnesha):

hmm nope 10 ??

OpenStudy (anonymous):

3*2 = 6 and 5*2 = 10

Nnesha (nnesha):

 multiply numerator by numerator and denominator by denominator